# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

### v2

@@ Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of agents. While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings. SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots. The code is publicly available at https://github.com/smart-mapf/smart.
Index Terms:
Multi-Agent Path Finding, Multi-Robot System, Realistic Simulation.
@@ I Introduction
The Multi-Agent Path Finding (MAPF) problem [1] focuses on planning collision-free paths for a large group of robots within a known environment. This problem finds various real-world applications [2, 3, 4], with warehouse automation being a prominent example, where hundreds of robots must be coordinated to perform daily operations. In recent years, MAPF has grown into a rapidly expanding research area [5, 6]. State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds [7, 8, 9]. However, these algorithms make several simplifying assumptions. First, they rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Therefore, while researchers have shown that MAPF algorithms have a wide range of applications, their performance in more realistic settings is unclear.
@@
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort. While some software tools [10, 11] have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
@@
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios. SMART comprises three main components: (1) physics-engine-based simulators, (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark [1]. The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) [3], a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and temporal uncertainty during execution.
@@
Our contributions in this paper are as follows: 1. We present SMART, an open-source software tool for evaluating MAPF algorithms in realistic scenarios. SMART directly accepts MAPF plans and problem instances from existing benchmarks. 2. We integrate SMART with two simulators, ARGoS3 and Isaac Sim, and robot models and empirically evaluate its scalability and replicability. SMART can scale to thousands of robots while maintaining consistent and reproducible results, while prior tools fail to handle a hundred robots. 3. We develop an online interface with interactive visualization, allowing users to easily monitor execution progress and access detailed statistical data during simulation. 4. We validate SMART on real-world mobile robots, demonstrating its robustness and effectiveness in physical environments.
@@
Figure 2: (a) MAPF problem. (b) Associated MAPF plan. (c) ADG from the MAPF plan. Green ovals are actions, with the first element as the reference time from the plan, the second as the action type, and the last two as start and goal locations. Horizontal edges are Type-1, and vertical edges are Type-2. All robots initially face north.
@@
II Background
In this section, we first introduce the MAPF problem and related work. Next, we discuss existing execution frameworks for executing MAPF plans. Finally, we review benchmarks used to evaluate MAPF algorithms.
II-A MAPF Problem and MAPF Plan
A MAPF instance [1] consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ constructed from a 2-D grid map and a set of $I$ robots $\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph are shown in Fig. 2 (a) and (b), respectively. The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
Prior research has proposed many MAPF algorithms, most of which are developed based on 2-D grid maps. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees [12, 13], while others focus on scalability, solving MAPF problems involving thousands of robots [14, 15]. Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency [16, 17]. Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process [18, 19, 20, 21]. They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
II-B Executing MAPF Plans
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging [22], especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the uncertainty in execution time. One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred [22]. The state-of-the-art execution framework is Action Dependency Graph (ADG) [3], which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various robot models [23].
II-C Evaluating MAPF Plans
To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF benchmarks with 2-D-grid maps are proposed [1, 24, 25, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space. Some benchmarks have attempted to include realistic elements such as acceleration and rotation [27, 28]. However, their model is still far from realistic, and typically requires the same robot models for both planning and execution phases.
Given the difficulty of modeling all real-world factors analytically, physics-engine-based simulators offer a promising approach for more accurate evaluation. MRP-Bench [10] leverages a high-fidelity, physics-based simulator Gazebo [29] to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC [11] also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, these existing tools have two main limitations. First, they often require significant engineering effort to integrate user-defined maps and MAPF planners, limiting their flexibility for different scenarios. In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners using different robot models. Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 agents. Most of these works rely on Gazebo, whose high computational overhead, resulting from modeling many factors not relevant to MAPF, significantly constrains scalability. MRP-Bench, for example, does not scale to more than 10 robots.
@@ III-A Simulator
Given a user-provided configuration, the simulator creates realistic simulation environments. The configuration specifies obstacle locations and shapes, robot start locations, and detailed robot configurations (e.g., the robot model, as well as its velocity and acceleration limits). To support the widely used MovingAI benchmark [1], SMART can converts its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator. During execution, the simulator uses a physics engine to update the states of the robots and environment, including their poses, dynamics, sensors, and actuators, based on the received control commands from the executors. The executors can query the simulator for real-time robot information, such as their current pose. To accommodate diverse user requirements, SMART supports two simulation platforms: ARGoS3 [30] and Isaac Sim [31].
@@ III-A2 Isaac Sim
Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic.
@@ III-B Execution Monitoring (EM) Server
During the initialization, we use the MAPF plan parser to construct an ADG. We define a standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$. The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans. This allows SMART to work with plans using varying simplified robot models. The parser first converts the path of each robot into a sequence of actions. Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path and adds a rotation action when needed. The ADG is constructed based on the sequence of actions. ADG requires the MAPF plan to be circle-conflict free, ensuring that robots do not collide even if delayed by one timestep [3]. However, ensuring this is not standard in most MAPF models. As a result, if we directly build an ADG based on a plan with circle conflict, we may create deadlocks, such as the one illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by [32], we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
Execution Managing
During execution, an execution monitor framework manages the progress of execution and communicates with robot executors. This framework follows the method described in ADG [3] to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to enqueued if (1) it has no preceding vertices, or (2) its preceding vertex connected by a Type-1 edge is enqueued or finished, and all its preceding vertices connected by Type-2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with robot executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
@@ III-C Executor
Each robot is launched with an individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client establishes communication with the server, receiving enqueued actions and sending confirmation for finished actions. The action queue keeps track of the enqueued actions, which are executed in a first-in-first-out manner. A Proportional Integral Derivative (PID) controller is used to execute these actions by sending control commands to the simulator. Once an action is finished, a confirmation message is sent to the server. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue. For example, if a robot receives three consecutive forward movement actions, the executor merges them into a single longer forward action, reducing unnecessary stop-and-go behavior.

### v3

# Advancing MAPF Toward the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

@@ Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of robots. While state-of-the-art MAPF planners can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF planners in their specific settings. SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF planners and robot models, and (3) SMART scales to thousands of robots. The code is publicly available at https://github.com/smart-mapf/smart with an online service available at https://smart-mapf.github.io/demo/.
@@ I Introduction
The Multi-Agent Path Finding (MAPF) problem [29] focuses on planning collision-free paths for a large group of robots within a known environment. This problem finds various real-world applications [18, 11, 10], with warehouse automation being a prominent example, where hundreds of robots must be coordinated to perform daily operations. In recent years, MAPF has grown into a rapidly expanding research area [19, 32]. State-of-the-art MAPF planners can plan paths for hundreds of robots within seconds [16, 35, 14]. However, these algorithms make several simplifying assumptions. First, they rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Therefore, while researchers have shown that MAPF planners have a wide range of applications, their performance in more realistic settings is unclear.
@@
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF planners in realistic settings. In this paper, we use the term scalability to refer to runtime capacity in terms of the maximum number of robots that the testbed can support in simulation. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF planners is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such robotic simulation tools can require significant effort. While some software tools [27, 9] have been developed to evaluate MAPF planners in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF planners and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF planners in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF planners, industrial professionals can more easily understand and select a MAPF planner that fits their demands.
@@
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF planners in realistic scenarios. SMART comprises three main components: (1) a physics-engine-based simulator, (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark [29]. The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) [11], a MAPF execution framework, to ensure robust execution of the paths generated by MAPF planners using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and temporal uncertainty during execution.
@@
Our contributions in this paper are as follows: 1. We present SMART, an open-source software tool for evaluating MAPF planners in realistic scenarios. SMART directly accepts MAPF plans and problem instances from existing benchmarks. 2. We integrate SMART with two simulators, ARGoS3 [25] and Isaac Sim [21], and empirically evaluate its scalability and replicability. SMART can scale to thousands of robots while maintaining consistent and reproducible results, whereas prior tools fail to handle a hundred robots. 3. We develop an online interface with interactive visualization, allowing users to easily monitor execution progress and access detailed statistical data during simulation. 4. We validate SMART on real-world mobile robots, demonstrating its robustness and effectiveness in physical environments.
@@
Figure 2: (a) MAPF problem. (b) Associated MAPF plan. (c) ADG from the MAPF plan. Green ovals are actions, with the first element as the reference time from the plan, the second as the action type (e.g., rotation or forward movement), and the last two as start and goal locations. Horizontal edges are Type-1, and vertical edges are Type-2. All robots initially face north.
II Background
In this section, we first introduce the MAPF problem and related work. Next, we discuss existing execution frameworks for executing MAPF plans. Finally, we review benchmarks used to evaluate MAPF planners.
II-A MAPF Problem and MAPF Plan
A MAPF instance [29] consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ constructed from a 2-D grid map and a set of $I$ robots $\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF planner aims to find a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph are shown in Fig. 2 (a) and (b), respectively. The typical objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
Prior research has proposed many MAPF planners, most of which were implemented and tested on 2-D grid maps. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees [28, 15], while others focus on scalability, solving MAPF problems involving thousands of robots [22, 23]. Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process [5, 1, 33, 34]. They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among robots, while the low level plans for each robot considering real-world factors.
II-B Executing MAPF Plans
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging [3], especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the uncertainty in execution time. One naive way is synchronized execution, where robots need to wait for all robots to finish their current actions before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally expensive and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred [3]. The state-of-the-art execution framework is Action Dependency Graph (ADG) [11], which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various robot models [31].
Although modern MAPF solvers have become significantly faster, execution frameworks such as ADG remain essential due to inevitable real-world uncertainties. Some of them, such as communication delays, cannot be eliminated through faster planning alone. Moreover, by enforcing only necessary ordering constraints rather than strict timestep synchronization, ADG enables robots to execute multiple consecutive actions without stopping, supporting velocity optimization and smoother execution rather than requiring stop-and-go behavior at every timestep.
TABLE I: Comparison of SMART with existing multi-robot frameworks and evaluation tools. Interactive execution diagnostics refer to inspection of execution states during runtime, and benchmark support refers to compatibility with standard MAPF benchmarks.
Tool Max Robots MAPF Planner General Execution Interactive Execution Benchmark Primary
(Tested) Integration Framework Diagnostics Support Purpose
Open-RMF [24] $\sim$100 Manual Yes No No Fleet deployment
MRP-Bench [27] <<10 Manual No No Partial Small-scale MAPF evaluation
REMROC [9] <<50 Manual No No No Human-aware evaluation
SMART (ours) 2000 Yes Yes Yes Yes MAPF research
II-C Evaluating MAPF Plans
To evaluate MAPF planners, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF benchmarks with 2-D grid maps are proposed [29, 8, 30, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space. Some benchmarks have attempted to include realistic elements such as acceleration and rotation [20, 4]. However, their models are still far from realistic, and they typically require the same robot models for planning and execution phases.
Given the difficulty of modeling all real-world factors analytically, physics-engine-based simulators offer a promising approach for more accurate evaluation. Frameworks like Open-RMF [24] provide middleware for deploying and coordinating robot fleets in production environments, handling task allocation, traffic management, and building infrastructure integration. However, such operational deployment frameworks are not specialized for MAPF evaluation. While centralized MAPF planners can be integrated, using them to benchmark and stress-test MAPF planners at large scale typically requires additional engineering effort beyond their core scope. To address the need for MAPF-specific evaluation, several simulation-based tools have been developed. MRP-Bench [27] leverages a high-fidelity, physics-based simulator Gazebo [13] to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC [9] also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, as shown in Table I, these existing tools have two main limitations. First, they often require significant engineering effort to integrate user-defined maps and new MAPF planners, limiting their flexibility for different scenarios. In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners across different robot models assumed during planning. Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 robots. Most of these works rely on Gazebo, whose high computational overhead, resulting from modeling many factors not relevant to MAPF, significantly constrains scalability. MRP-Bench, for example, does not scale to more than 10 robots.
@@ III-A Simulator
Given a user-provided configuration, the simulator creates a realistic simulation environment. The configuration specifies obstacle locations and shapes, robot start locations, and robot configurations (e.g., velocity and acceleration limits). To support the widely used MovingAI benchmark [29], SMART can convert its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator. During execution, the simulator uses a physics engine to update the states of the robots, including their poses, dynamics, and actuators, based on the received control commands from the executors. The simulator provides real-time robot state information (e.g., current pose) to other system components. To accommodate diverse user requirements and balance scalability with physical realism, SMART supports two simulation platforms: ARGoS3 [25] for efficient large-scale experiments and Isaac Sim [21] for higher-fidelity simulation.
@@ III-A2 Isaac Sim
Figure 4: (a) Cycle in the ADG causing all robots to be stalled. (b) Adding safe stops between vertices to make the ADG acyclic. Robot diameter must be at most half the cell size to prevent collisions at safe stops between vertices.
@@ III-B Execution Monitoring (EM) Server
During initialization, we use the MAPF plan parser to convert it into a sequence of actions and then construct an ADG based on these actions. We define a lightweight, planner-agnostic standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$. This simple representation can be easily derived from any MAPF-generated path, making the parser compatible with any valid MAPF plan regardless of the underlying robot model or planner used to generate them. This design choice allows SMART to work seamlessly with planners using varying simplified robot models without requiring modifications to the planners themselves. The parser first converts the path of each robot into a sequence of actions. Specifically, for each transition between consecutive vertices, the parser infers the required robot orientation. If the orientation differs from the current orientation, an in-place rotation action is added. Subsequently, a forward movement action is added to move the robot to the next vertex. The ADG is then constructed based on this sequence of actions. ADG requires the MAPF plan to be cycle-conflict-free, i.e., it prohibits cyclic location exchanges among a set of robots within the same timestep. This is because ADG only allows movements into empty space to ensure delay safety, which cannot be satisfied under a cycle conflict [11]. However, ensuring this is not standard in most MAPF models. As a result, if we directly build an ADG based on a plan with cycle conflicts, we may create deadlocks, such as the one illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by [12], we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
Execution Management
During execution, an execution monitor framework manages the progress of execution and communicates with robot executors. This framework follows the method described in ADG [11] to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to enqueued if (1) it has no preceding vertices, or (2) its preceding vertex connected by a Type-1 edge is enqueued or finished, and all its preceding vertices connected by Type-2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with executors is managed by a Remote Procedure Call (RPC) server built with rpclib. [^1] Available at http://rpclib.net.
@@ III-C Executor
Each robot is launched with an individual executor. As shown in Fig. 3, each executor consists of an RPC client, an action queue, and a controller. The RPC client communicates with the server by receiving enqueued actions and sending confirmations for finished actions. The action queue stores enqueued actions and executes them in first-in-first-out order. A Proportional Integral Derivative (PID) controller executes these actions by sending control commands to the simulator, and a confirmation is sent to the server upon completion. To allow robots to operate at higher speeds, consecutive actions of the same type are merged into one action in the queue. For example, if a robot receives three consecutive forward movement actions, the executor merges them into a single longer forward action, reducing unnecessary stop-and-go behavior. This distributed executor design enables parallel operation, as each robot handles the above-mentioned operations independently, improving scalability.